OpenStudy (anonymous):

A bag contains 4 black marbles and 6 white marbles. 3 marbles are selected randomly from the bag. Find the probability of selecting 3 white marbles with replacement

OpenStudy (kira_yamato):

P(WWW) = (6/10)^3 = 27/125